### v1

# Learnability in Online Kernel Selection with Memory Constraint via Data-dependent Regret Analysis

Abstract
Online kernel selection is a fundamental problem of online kernel methods. In this paper, we study online kernel selection with memory constraint in which the memory of kernel selection and online prediction procedures is limited to a fixed budget. An essential question is what is the intrinsic relationship among online learnability, memory constraint and data complexity? To answer the question, it is necessary to show the trade-offs between regret bound and memory constraint. Previous work gives a worst-case lower bound depending on the data size, and shows learning is impossible within a small memory constraint. In contrast, we present a different result by providing data-dependent upper bounds depending on two data complexities, namely kernel alignment and the cumulative losses of competitive hypothesis. We propose an algorithmic framework giving data-dependent upper bounds for two types of loss functions. For the hinge loss function, our algorithm achieves an expected upper bound depending on kernel alignment. For smooth loss functions, our algorithm achieves a high-probability upper bound depending on the cumulative losses of competitive hypothesis. We also prove a matching lower bound for smooth loss functions. Our results show that if the two data complexities are sub-linear, then learning is possible within a small memory constraint. Our algorithmic framework depends on a new buffer maintaining framework and a reduction from online kernel selection to prediction with expert advice. Finally, we empirically verify the prediction performance of our algorithms on benchmark datasets.
1 Introduction
Online kernel selection (OKS) aims to dynamically select a kernel function or a reproducing kernel Hilbert space (RKHS) for online kernel learning algorithms. Compared with offline kernel selection, OKS poses more computational challenges for online kernel selection algorithms due to both the kernel selection procedure and prediction needing to be real-time. We can formulate it as a sequential decision problem. Let $\mathcal{K}=\{\kappa_{i}\}^{K}_{i=1}$ contain $K$ base kernel functions and $\mathcal{H}_{i}$ be a dense subset of the RKHS induced by $\kappa_{i}$. At round $t$, an adversary selects an instances ${\bm{x}}_{t}\in\mathbb{R}^{d}$. Then a learner chooses a hypotheses $f_{t}\in\cup^{K}_{i=1}\mathcal{H}_{i}$ and makes a prediction. The adversary gives the true output $y_{t}$, and the learner suffers a loss $\ell(f_{t}({\bm{x}}_{t}),y_{t})$. We use the regret to measure the performance of the learner which defined as follows,
$\forall\kappa_{i}\in\mathcal{K},\quad\forall f\in\mathcal{H}_{i},~{}\mathrm{Reg}(f)=\sum^{T}_{t=1}\ell(f_{t}({\bm{x}}_{t}),y_{t})-\sum^{T}_{t=1}\ell(f({\bm{x}}_{t}),y_{t}).$
An efficient algorithm should guarantee $\mathrm{Reg}(f)=o(T)$. For convex loss functions, many algorithms [1, 2, 3] achieve (or imply)
$\mathrm{Reg}(f)=\tilde{O}\left(\left(\|f\|^{\alpha}_{\mathcal{H}_{i}}+1\right)\sqrt{T\ln{K}}\right),\alpha\in\{1,2\}.$
The algorithms can adapt to the optimal, yet unknown hypothesis space within a small information-theoretical cost.
A major challenge of OKS is the so called curse of dimensionality, i.e., algorithms must store the previous $t-1$ instances at the $t$-th round. The $O(t)$ memory cost is prohibitive for large-scale online learning problems. To address this issue, we limit the memory of algorithms to a budget of $\mathcal{R}$ quanta. For convex loss functions, the worst-case regret is $\Theta(\max\{\sqrt{T},\frac{T}{\sqrt{\mathcal{R}}}\})$ [4] which proves a trade-off between memory constraint and regret. To be specific, achieving a $O(T^{\alpha})$ regret requires $\mathcal{R}=\Omega(T^{2(1-\alpha)})$, $\alpha\in[\frac{1}{2},1)$. A constant or $\Theta(\ln{T})$ memory cost can not ensure sub-linear regret and thus learning is impossible in any $\mathcal{H}_{i}$. We use the regret to define the learnability of a hypothesis space (Definition 2). However, empirical results have shown that a small memory cost is enough to achieve good prediction performance [5, 6, 7]. It seems that learning is possible in the optimal hypothesis space. To fill the gap between learnability and regret, it is necessary to establish data-dependent regret bounds. To this end, we focus on the following two questions:
  1. Q1
    What trade-offs exist between memory constraint and certain data-dependent regret bounds, that is, how do data-dependent regret bounds depend on $\mathcal{R}$, $T$, and $K$?
  2. Q2
    Sub-linear regret implies that some hypothesis spaces are learnable. Is it possible to achieve sub-linear regret within a $O(\ln{T})$ memory cost?
In this paper, we will answer the two questions affirmatively. We first propose an algorithmic framework. Then we apply the algorithmic framework to two types of loss functions and achieve two kinds of data-dependent regret bounds. To maintain the memory constraint, we reduce it to the budget that limits the size of the buffer. Our algorithmic framework will use the buffer to store a subset of the observed examples. The main results are summarized as follows:
  1. 1)
    For the hinge loss function, our algorithm achieves an expected kernel alignment regret as follows (see Theorem 1 for detailed result),
    $\forall\kappa_{i}\in\mathcal{K},~{}\forall f\in\mathbb{H}_{i},~{}\mathbb{E}[\mathrm{Reg}(f)]=\tilde{O}\left(\sqrt{L_{T}(f)\ln{K}}+\frac{\sqrt{\mathcal{R}}}{\sqrt{K}}+\frac{\sqrt{K}}{\sqrt{\mathcal{R}}}\mathcal{A}_{T,\kappa_{i}}\right),$ (1)
    where $\mathcal{A}_{T,\kappa_{i}}$ is called kernel alignment, $\mathbb{H}_{i}\subseteq\mathcal{H}_{i}$ and
    $L_{T}(f)=\sum^{T}_{t=1}\ell(f({\bm{x}}_{t}),y_{t}).$ (2)
  2. 2)
    For smooth loss functions (Assumption 3), our algorithm achieves a high-probability small-loss bound (see Theorem 3 for detailed result). $\forall\kappa_{i}\in\mathcal{K},\forall f\in\mathbb{H}_{i}$, with probability at least $1-\delta$,
    $\begin{split}\mathrm{Reg}(f)=\tilde{O}\left(\frac{L_{T}(f)}{\sqrt{\mathcal{R}}}+\sqrt{L_{T}(f)\ln{K}}+\sqrt{\mathcal{R}}\right).\end{split}$ (3)
  3. 3)
    For smooth loss functions, we prove a lower bound on the regret (see Theorem 4 for detailed result). Let $K=1$. The regret of any algorithm that stores $B$ examples satisfies
    $\exists f\in\mathbb{H},\quad\mathbb{E}\left[\mathrm{Reg}(f)\right]=\Omega\left(U\cdot\frac{L_{T}(f)}{\sqrt{B}}\right),$
    where $U>0$ is a constant and bounds the norm of all hypotheses in $\mathbb{H}$. The upper bound in (3) is optimal in terms of the dependence on $L_{T}(f)$.
The data-dependent regret bounds in (1) and (3) improve the $O(\sqrt{T\ln{K}}+\|f\|^{2}_{\mathcal{H}_{i}}\max\{\sqrt{T},\frac{T}{\sqrt{\mathcal{R}}}\})$ [4] bound, since $\mathcal{A}_{T,\kappa_{i}}=O(T)$ and $L_{T}(f)=O(T)$. If $\kappa_{i}$ matches well with the data, then we expect $\mathcal{A}_{T,\kappa_{i}}\ll T$ or $L_{T}(f)\ll T$. In the worst case, i.e., $\mathcal{A}_{T,\kappa_{i}}=\Theta(T)$ or $L_{T}(f)=\Theta(T)$, our bounds are same with previous result. We give new trade-offs between memory constraint and regret bound which answers Q1. For online kernel selection, we just aim to adapt to the optimal kernel $\kappa^{\ast}\in\mathcal{K}$. If $\kappa^{\ast}$ matches well with the data, then we expect $\mathcal{A}_{T,\kappa^{\ast}}=o(T)$ and $\min_{f\in\mathcal{H}_{\kappa^{\ast}}}L_{T}(f)=o(T)$. In this case, a $\Theta(\ln{T})$ memory cost is enough to achieve sub-linear regret which answers Q2. Thus learning is possible in $\mathbb{H}_{\kappa^{\ast}}$ within a small memory constraint.
Our algorithmic framework reduces OKS to prediction with expert advice and uses the well-known optimistic mirror descent framework [8, 9]. We also propose a new buffer maintaining framework.
Related Work
Previous work used buffer maintaining technique and random features to develop algorithms for online kernel selection with memory constraint [1, 3, 7, 10]. However, all of existing results can not answer Q1 and Q2 well. The sketch-based online kernel selection algorithm [7] enjoys a $O(B\ln{T})$ regularized regret (the loss function incorporates a regularizer) within a buffer of size $B$. The regret bound becomes $O(T\ln{T})$ in the case of $B=\Theta(T)$. The Raker algorithm [10] achieves a $O(\sqrt{T\ln{K}}+T/\sqrt{D})$  [^1] The original regret bound is $O(U\sqrt{T\ln{K}}+\epsilon TU)$, and holds with probability $1-\Theta\left(\epsilon^{-2}\exp(-D\epsilon^{2}/(4d+8))\right)$. regret and suffers a space complexity in $O((d+K)D)$, where $D$ is the number of random features. The upper bound also shows a trade-off in the worst case. If $D$ is a constant, then Raker can not achieve a sub-linear regret bound which gives a negative answer to Q2.
Our work is also related to achieving data-dependent regret bounds for online learning and especially online kernel learning. For online learning, various data-dependent regret bounds including small-loss bound [11, 12, 13], variance bound [14, 15] and path-length bound [8, 16, 17], have been achieved. The adaptive online learning framework [18] can achieve data-dependent and model-dependent regret bounds, but can not induce a computationally efficient algorithm. For online kernel learning, achieving data-dependent regret bounds is harder, since we must balance the computational cost. For loss functions satisfying specific smoothness (Assumption 3), the OSKL algorithm [19] achieves a small-loss bound. For loss functions with a curvature property, the PROS-N-KONS algorithm [20] achieves a regret bound depending on the effective dimension of the kernel matrix. All of these algorithms did not consider memory constraint, and thus did not give a trade-off between memory constraint and regret bound.
2 Problem Setup
Let $\mathcal{I}_{T}:=\{({\bm{x}}_{t},y_{t})\}_{t\in[T]}$ be a sequence of examples, where ${\bm{x}}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d},y_{t}\in[-1,1]$ and $[T]:=\{1,\ldots,T\}$. Let $\kappa(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a positive semidefinite kernel function and $\mathcal{H}_{\kappa}$ be a dense subset of the associated RKHS, such that, for any $f\in\mathcal{H}_{\kappa}$,
  1. (i)
    $\langle f,\kappa({\bm{x}},\cdot)\rangle_{\mathcal{H}_{\kappa}}=f({\bm{x}}),\forall{\bm{x}}\in\mathcal{X}$,
  2. (ii)
    $\mathcal{H}_{\kappa}=\overline{\mathrm{span}(\kappa({\bm{x}}_{t},\cdot)|t\in[T])}$.
We define $\langle\cdot,\cdot\rangle_{\mathcal{H}_{\kappa}}$ as the inner product in $\mathcal{H}_{\kappa}$, which induces the norm $\|f\|_{\mathcal{H}_{\kappa}}=\sqrt{\langle f,f\rangle_{\mathcal{H}_{\kappa}}}$. For simplicity, we will omit the subscript $\mathcal{H}_{\kappa}$ in the inner product. Let (,):×[1$\ell(\cdot,\cdot):\mathbb{R}\times[-1,1]\rightarrow\mathbb{R}$ be the loss function. Denote by
$\mathcal{B}_{\psi}(f,g)=\psi(f)-\psi(g)-\langle\nabla\psi(g),f-g\rangle,~{}\forall f,g\in\mathcal{H}_{\kappa},$
the Bregman divergence associated with a strongly convex regularizer $\psi:\mathcal{H}_{\kappa}\rightarrow\mathbb{R}$.
If an oracle gives the optimal kernel $\kappa^{\ast}\in\mathcal{K}$ for $\mathcal{I}_{T}$, then we just run an online kernel learning algorithm in $\mathcal{H}_{\kappa^{\ast}}$. Since $\kappa^{\ast}$ is unknown, the learner aims to develop a kernel selection algorithm and generate a sequence of hypotheses $\{f_{t}\}^{T}_{t=1}$ which is competitive to any $f\in\mathcal{H}_{\kappa^{\ast}}$, that is, the learner hopes to guarantee $\mathrm{Reg}(f)=o(T)$. The representer theorem proves that $f=\sum^{T}_{t=1}a_{t}\kappa_{i}({\bm{x}}_{t},\cdot)$ for any $f\in\mathcal{H}_{i}$. Thus storing $f_{t}$ needs to store $\{({\bm{x}}_{\tau},y_{\tau})\}^{t-1}_{\tau=1}$ and leads to a memory cost in $O(t)$. In this paper, we consider online kernel selection with memory constraint. Next we define the memory constraint and reduces the memory constraint to the size of a buffer.
Definition 1 (Memory Budget [4]).
A memory budget of $\mathcal{R}$ quanta is the maximal memory that any online kernel selection algorithm can use.
Assumption 1 ([4]).
$\forall\kappa_{i}\in\mathcal{K}$, there is a constant $\alpha>0$, such that any budgeted online kernel leaning algorithm running in $\mathcal{H}_{i}$ can maintain a buffer of size $B\leq\alpha\mathcal{R}$ within $\mathcal{R}$ quanta memory constraint. If the space complexity of algorithm is linear with $B$, then “$=$” holds.
Assumption 1 implies that there is no need to assume $\mathcal{R}=\infty$ unless $T=\infty$, since the maximal $\mathcal{R}$ satisfies $B=T$. We define budgeted online kernel learning algorithms as such algorithms that operate on a subset of the observed examples, such as Forgetron [5], BSGD [21] to name but a few. In Assumption 1, $\alpha$ is independent of kernel function, since the memory cost is used to store the examples and the coefficients.
Assumption 2.
There is a positive constant $k_{1}$, such that $\forall\kappa_{i}\in\mathcal{K}$ and $\forall{\bm{u}}\in\mathcal{X}$, $\kappa_{i}({\bm{u}},{\bm{u}})\in[k_{1},1]$.
We will define two types of data complexities and use them to bound the regret. The first one is called kernel alignment which is defined as follows
$\mathcal{A}_{T,\kappa_{i}}=\sum^{T}_{t=1}\kappa_{i}({\bm{x}}_{t},{\bm{x}}_{t})-\frac{1}{T}{\bm{Y}}^{\mathrm{T}}{\bm{K}}_{i}{\bm{Y}},$
where ${\bm{Y}}=(y_{1},\ldots,y_{T})^{\mathrm{T}}$ and ${\bm{K}}_{i}$ is the kernel matrix induced by $\kappa_{i}$. The measure $\mathcal{A}_{T,\kappa_{i}}$ quantifies the matching between ${\bm{K}}_{i}$ and the examples. If ${\bm{K}}_{i}={\bm{Y}}{\bm{Y}}^{\mathrm{T}}$ is the ideal kernel matrix, then $\mathcal{A}_{T,\kappa_{i}}=\Theta(1)$. The second one is called small-loss defined by $\min_{f\in\mathbb{H}_{i}}L_{T}(f)$ where $L_{T}(f)$ follows (2) and
$\mathbb{H}_{i}=\left\{f\in\mathcal{H}_{i}:\|f\|_{\mathcal{H}_{i}}\leq U\right\}.$
Both $\mathcal{A}_{T,\kappa_{i}}$ and $L_{T}(f)$ are independent of algorithms.
Finally, we define the online learnability which is a variant of the learnability defined in [22].
Definition 2 (Online Learnability).
Given $\mathcal{I}_{T}$, a hypothesis space $\mathcal{H}$ is said to be online learnable if $\lim_{T\rightarrow\infty}\frac{\mathrm{Reg}(f)}{T}=0$, $\forall f\in\mathcal{H}$.
The definition is equivalent to $\mathrm{Reg}(f)=o(T)$. We can also replace $\mathrm{Reg}(f)$ with $\mathbb{E}[\mathrm{Reg}(f)]$. The learnability defined in [22] considers the whole sample space, i.e., the worst-case examples, while our definition only considers a fixed $\mathcal{I}_{T}$. In the worst-case, the two kinds of learnability are equivalent. Since $\mathcal{I}_{T}$ may not always be generated in the worst-case, our learnability can adapt to the hardness of $\mathcal{I}_{T}$.
3 Algorithmic Framework
We reduce OKS to prediction with expert advice (PEA) where $\kappa_{i}$ corresponds to the $i$-th expert. The main idea of our algorithmic framework is summarized as follows: (i) generating $\{f_{t,i}\in\mathcal{H}_{i}\}^{T}_{t=1}$ for all $i\in[K]$; (ii) aggregating the predictions $\{f_{t,i}({\bm{x}}_{t})\}^{K}_{i=1}$. The challenge is how to control the size of $\{f_{t,i}\}^{K}_{i=1}$. To this end, we propose a new buffer maintaining framework containing an adaptive sampling and a removing process.
3.1 Adaptive Sampling
We use the optimistic mirror descent framework (OMD) [8, 9] to generate $\{f_{t,i}\in\mathcal{H}_{i}\}^{T}_{t=1}$. OMD maintains two sequences of hypothesis, i.e., $\{f_{t,i}\}^{T}_{t=1}$ and $\{f^{\prime}_{t,i}\}^{T}_{t=1}$ which are defined as follows,
$f_{t,i}=$ $\mathop{\arg\min}_{f\in\mathcal{H}_{i}}\left\{\left\langle f,\hat{\nabla}_{t,i}\right\rangle+\mathcal{B}_{\psi_{i}}\left(f,f^{\prime}_{t-1,i}\right)\right\},$ (4)
$f^{\prime}_{t,i}=$ $\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\left\langle f,\nabla_{t,i}\right\rangle+\mathcal{B}_{\psi_{i}}\left(f,f^{\prime}_{t-1,i}\right)\right\},$ (5)
where $\nabla_{t,i}$ is the (sub)-gradient, ^$\hat{\nabla}_{t,i}$ is an optimistic estimator of $\nabla_{t,i}$ and
$\psi_{i}(f)=$ $\frac{1}{2\lambda_{i}}\|f\|^{2}_{\mathcal{H}_{i}},$
$\nabla_{t,i}=$ $\ell^{\prime}(f_{t,i}({\bm{x}}_{t}),y_{t})\cdot\kappa_{i}({\bm{x}}_{t},\cdot).$
Note that the hypothesis space in (4) and (5) is $\mathcal{H}_{i}$ and $\mathbb{H}_{i}$, respectively. This trick [23] is critical to obtaining regret bounds depending on kernel alignment. At the beginning of round $t$, we execute (4) and output $f_{t,i}({\bm{x}}_{t})$. The key is to control the size of $f^{\prime}_{t,i}$. Using Assumption 1, we construct a buffer denoted by $S_{i}$ which stores the examples constructing $f^{\prime}_{t,i}$. Thus we just need to limit the size of $S_{i}$. To this end, we propose an adaptive sampling scheme. Let $b_{t,i}$ be a Bernoulli random variable and
$\mathbb{P}[b_{t,i}=1]=\frac{\|\nabla_{t,i}-\hat{\nabla}_{t,i}\|^{\nu}_{\mathcal{H}_{i}}}{Z_{t}},\quad\nu\in\{1,2\},$ (6)
where $Z_{t}$ is a normalizing constant. We further define
$\mathrm{con}(a(i)):=$ $\left\|\kappa_{i}({\bm{x}}_{i(s_{t})},\cdot)-\kappa_{i}({\bm{x}}_{t},\cdot)\right\|_{\mathcal{H}_{i}}\leq\gamma_{t,i},$
${\bm{x}}_{i(s_{t})}=$ $\mathop{\arg\min}_{{\bm{x}}_{\tau}\in S_{i}}\left\|\kappa_{i}({\bm{x}}_{\tau},\cdot)-\kappa_{i}({\bm{x}}_{t},\cdot)\right\|_{\mathcal{H}_{i}}.$
$\mathrm{con}(a(i))$ means that if there is an instance in $S_{i}$ that is similar with ${\bm{x}}_{t}$, then we can use it as a proxy of ${\bm{x}}_{t}$. In this way, $({\bm{x}}_{t},y_{t})$ will not be added into $S_{i}$.
If $\nabla_{t,i}=0$, then $S_{i}$ keeps unchanged and we execute (5). If $\nabla_{t,i}\neq 0$ and $\mathrm{con}(a(i))$, then $S_{i}$ keeps unchanged and we replace (5) with (7).
$f^{\prime}_{t,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\left\langle f,\nabla_{i(s_{t}),i}\right\rangle+\mathcal{B}_{\psi_{i}}\left(f,f^{\prime}_{t-1,i}\right)\right\}.$ (7)
If $\nabla_{t,i}\neq 0,\neg\mathrm{con}(a(i))$, then we replace (5) with (8).
$f^{\prime}_{t,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\left\langle f,\tilde{\nabla}_{t,i}\right\rangle+\mathcal{B}_{\psi_{i}}\left(f,f^{\prime}_{t-1,i}\right)\right\}.$ (8)
In this case, if $b_{t,i}=1$, then we add $({\bm{x}}_{t},y_{t})$ into $S_{i}$. In the above two equations, we define
$\nabla_{i(s_{t}),i}=$ $\ell^{\prime}(f_{t,i}({\bm{x}}_{t}),y_{t})\cdot\kappa_{i}({\bm{x}}_{i(s_{t})},\cdot),$
~$\tilde{\nabla}_{t,i}=$ $\frac{\nabla_{t,i}-\hat{\nabla}_{t,i}}{\mathbb{P}[b_{t,i}=1]}\mathbb{I}_{b_{t,i}=1}+\hat{\nabla}_{t,i}.$
3.2 Removing a Half of Examples
We allocate a memory budget $\mathcal{R}_{i}$ for $f^{\prime}_{t,i}$ and require $\sum^{K}_{i=1}\mathcal{R}_{i}\leq\mathcal{R}$. Using Assumption 1, we just need to limit $|S_{i}|\leq\alpha\mathcal{R}_{i}$. At any round $t$, if $|S_{i}|=\alpha\mathcal{R}_{i}$ and $b_{t,i}=1$, then we remove a half of examples from $S_{i}$ [23]. Let $S_{i}=\{({\bm{x}}_{r_{j}},y_{r_{j}})\}^{\alpha\mathcal{R}_{i}}_{j=1}$. We rewrite $f^{\prime}_{t-1,i}$ as follows,
$f^{\prime}_{t-1,i}=$ $f^{\prime}_{t-1,i}(1)+f^{\prime}_{t-1,i}(2),$
$f^{\prime}_{t-1,i}(1)=$ $\sum^{\alpha\mathcal{R}_{i}/2}_{j=1}\beta_{r_{j}}\kappa({\bm{x}}_{r_{j}},\cdot),$
$f^{\prime}_{t-1,i}(2)=$ $\sum^{\alpha\mathcal{R}_{i}}_{j=\alpha\mathcal{R}_{i}/2+1}\beta_{r_{j}}\kappa({\bm{x}}_{r_{j}},\cdot).$
Let $o_{t}\in\{1,2\}$ and $q_{t}\in\{1,2\}\setminus\{o_{t}\}$. We will remove $f^{\prime}_{t-1,i}(o_{t})$ from $f^{\prime}_{t-1,i}$, that is, the examples constructing $f^{\prime}_{t-1,i}(o_{t})$ are removed from $S_{i}$. We further project $f^{\prime}_{t-1,i}(q_{t})$ onto $\mathbb{H}_{i}$ by
$\bar{f}^{\prime}_{t-1,i}(q_{t})=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\|f-f^{\prime}_{t-1,i}(q_{t})\right\|^{2}_{\mathcal{H}}.$
Then we execute the second mirror descent and redefine (8) as follows
$f^{\prime}_{t,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\left\langle f,\tilde{\nabla}_{t,i}\right\rangle+\mathcal{B}_{\psi_{i}}\left(f,\bar{f}^{\prime}_{t-1,i}(q_{t})\right)\right\}.$ (9)
Our buffer maintaining framework always ensures $|S_{i}|\leq\alpha\mathcal{R}_{i}$. A natural approach is the restart technique which removes all of the examples in $S_{i}$, i.e., resetting $S_{i}=\emptyset$. We will prove that removing a half of examples can achieve better regret bounds.
3.3 Reduction to PEA
Similar to previous online multi-kernel learning algorithms [1, 10, 24], we use PEA framework to aggregate the outputs of $K$ hypotheses. Let ${\bm{p}}_{t}\in\Delta_{K-1}$ be a probability distribution over $\{f_{t,i}\}^{K}_{i=1}$. At each round $t$, we output the prediction $f_{t}({\bm{x}}_{t})=\sum^{K}_{i=1}p_{t,i}f_{t,i}({\bm{x}}_{t})$ or $\hat{y}_{t}=\mathrm{sign}(f_{t}({\bm{x}}_{t})$. We compute ${\bm{p}}_{t+1}$ using the multiplicative-weight update (Subsection 2.1 in [11]),
$\begin{split}p_{t+1,i}&=\frac{w_{t+1,i}}{\sum^{K}_{j=1}w_{t+1,j}},\\ w_{t+1,i}&=\exp\left(-\eta_{t+1}\sum^{t}_{\tau=1}c_{\tau,i}\right),\\ \eta_{t+1}=&\frac{\sqrt{2\ln{K}}}{\sqrt{1+\sum^{t}_{\tau=1}\sum^{K}_{i=1}p_{\tau,i}c^{2}_{\tau,i}}},\eta_{1}=\sqrt{2\ln{K}},\end{split}$ (10)
where $c_{t,i}=g(f_{t,i}({\bm{x}}_{t}),y_{t})$, $g:\mathbb{R}^{2}\rightarrow\mathbb{R}^{+}\cup\{0\}$ and will be defined later.
4 Applications
In this section, we apply the proposed algorithmic framework to two types of loss functions and derive two kinds of data-dependent regret bounds.
4.1 Hinge Loss Function
We consider online binary classification tasks. Let the loss function be the Hinge loss, implying
$\nabla_{t,i}=-y_{t}\kappa_{i}({\bm{x}}_{t},\cdot)\cdot\mathbb{I}_{y_{t}f_{t,i}({\bm{x}}_{t})<1}.$
We use the “Reservoir Sampling (RS)” technique [14, 25] to create ^$\hat{\nabla}_{t,i}$. To this end, we create a new buffer $V$ with size $M\geq 1$. At the beginning of round $t$, we define
^$\hat{\nabla}_{t,i}=-\frac{1}{|V|}\sum_{({\bm{x}},y)\in V}y\cdot\kappa_{i}({\bm{x}},\cdot).$
RS is defined follows. At any round $t$, we add $({\bm{x}}_{t},y_{t})$ into $V$ with probability $\min\{1,\frac{M}{t}\}$. If $|V|=M$ and we will add the current example, then an old example must be removed from $V$ uniformly. We create another buffer $S_{0}$ that stores all of the examples added into $V$. It is easy to prove that $\mathbb{E}[|S_{0}|]\leq M(1+\ln{T})$. Let $\mathcal{R}_{0}=\frac{M(1+\ln{T})}{\alpha}$ and $\mathcal{R}_{i}=\frac{\mathcal{R}-\mathcal{R}_{0}}{K}$ where we require $\mathcal{R}\geq 2\mathcal{R}_{0}$. It is easy to verify that $\sum^{K}_{i=0}\mathcal{R}_{i}=\mathcal{R}$.
We instantiate the sampling scheme (6) as follows,
$\nu=2,\quad Z_{t}=\left\|\nabla_{t,i}-\hat{\nabla}_{t,i}\right\|^{2}_{\mathcal{H}_{i}}+\left\|\hat{\nabla}_{t,i}\right\|^{2}_{\mathcal{H}_{i}}.$
The additional ^$\|\hat{\nabla}_{t,i}\|^{2}_{\mathcal{H}_{i}}$ in $Z_{t}$ makes $\mathbb{P}[b_{t,i}=1]<1$ and is important for controlling the times of removing operation, while it also increases the regret. The trick that uses different hypothesis spaces in (4) and (5), gives a negative term in the regret bound. The negative term can cancel out ^$\|\hat{\nabla}_{t,i}\|^{2}_{\mathcal{H}_{i}}$. Let $o_{t}=2$. We remove $f^{\prime}_{t-1,i}(2)$ from $f_{t,i}$. Then (9) becomes
$f^{\prime}_{t,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\left\langle f,\tilde{\nabla}_{t,i}\right\rangle+\mathcal{B}_{\psi_{i}}\left(f,\bar{f}^{\prime}_{t-1,i}(1)\right)\right\}.$ (11)
At the $t$-th round, our algorithm outputs the prediction $\hat{y}_{t}=\mathrm{sign}(f_{t}({\bm{x}}_{t}))$. Let ${\bm{p}}_{1}$ be the uniform distribution and ${\bm{p}}_{t+1}$ follow (10) where $c_{t,i}=\ell(f_{t,i}({\bm{x}}_{t}),y_{t})$. We name this algorithm M-OMD-H (Memory bounded OMD for the Hinge loss) and present the pseudo-code in Algorithm 1.
Algorithm 1 M-OMD-H
0:  $\{\lambda_{i}\}^{K}_{i=1}$, $\{\eta_{t}\}^{T}_{t=1}$, $\mathcal{R}$, $U$.
0:  $f^{\prime}_{0,i}=\hat{\nabla}_{1,i}=0$, $S_{i}=V=\emptyset$.
1:  for $t=1,2,\ldots,T$ do
2:     Receive ${\bm{x}}_{t}$
3:     for $i=1,\ldots,K$ do
4:        Find ${\bm{x}}_{i(s_{t})}$
5:        Compute ^$\hat{\nabla}_{t,i}=\frac{-1}{|V|}\sum_{({\bm{x}},y)\in V}y\kappa_{i}({\bm{x}},\cdot)$
6:        Compute $f_{t,i}({\bm{x}}_{t})$ according to (4)
7:     end for
8:     Output $\hat{y}_{t}=\mathrm{sign}(f_{t}({\bm{x}}_{t}))$ and receive $y_{t}$
9:     for $i=1,\ldots,K$ do
10:        if $y_{t}f_{t,i}({\bm{x}}_{t})<1$ then
11:           if $\mathrm{con}(a(i))$ then
12:              Update $f^{\prime}_{t,i}$ following (7)
13:           else
14:              Sample $b_{t,i}\sim\mathrm{Ber}(\mathbb{P}[b_{t,i}=1],1)$
15:              if $b_{t,i}=1$ and $|S_{i}|\leq\alpha\mathcal{R}_{i}-1$ then
16:                 Update $f^{\prime}_{t,i}$ following (8)
17:              end if
18:              if $b_{t,i}=1$ $\mathrm{and}$ $|S_{i}|=\alpha\mathcal{R}_{i}$ then
19:                 Compute $\bar{f}^{\prime}_{t-1,i}(1)$
20:                 Update $f^{\prime}_{t,i}$ following (11)
21:                 Remove the latest $\frac{\alpha\mathcal{R}_{i}}{2}$ examples from $S_{i}$
22:              end if
23:              Update $S_{i}=S_{i}\cup\{({\bm{x}}_{t},y_{t}):\mathbb{I}_{b_{t,i}=1}\}$
24:           end if
25:        end if
26:        Compute ${\bm{p}}_{t+1}$ by (10)
27:        Update reservoir $V$ and $S_{0}$
28:     end for
29:  end for
Lemma 1.
Let $M>1$ and $\alpha\mathcal{R}:=B\geq 2M(1+\ln{T})$. For any $\mathcal{I}_{T}$, the expected times that M-OMD-H executes removing operation on $S_{i}$ are $\left\lceil\frac{12K}{Bk_{1}}\mathcal{A}_{T,\kappa_{i}}(1+\frac{\ln{T}}{M})\right\rceil$ at most.
Note that that the times of removing operation depend on the kernel alignment, not the time horizon $T$.
Theorem 1.
Let $\mathcal{R}$ satisfy Lemma 1 and $B>K$. Let $U>1$, $\lambda_{i}=U\frac{\sqrt{K}}{\sqrt{2B}}$, $\eta_{t}$ follow (10) and
$\gamma_{t,i}=\frac{\left\|\nabla_{t,i}-\hat{\nabla}_{t,i}\right\|^{2}_{\mathcal{H}_{i}}}{\sqrt{1+\sum_{\tau\leq t}\left\|\nabla_{\tau,i}-\hat{\nabla}_{\tau,i}\right\|^{2}_{\mathcal{H}_{i}}\cdot\mathbb{I}_{\nabla_{\tau,i}\neq 0}}}.$ (12)
The expected regret of M-OMD-H satisfies,
$\forall\kappa_{i}\in\mathcal{K},f\in\mathbb{H}_{i},\quad\mathbb{E}\left[\mathrm{Reg}(f)\right]=O\left(\sqrt{UL_{T}(f)\ln{K}}+\frac{U\sqrt{B}}{\sqrt{K}}+\frac{\sqrt{K}U\mathcal{A}_{T,\kappa_{i}}\ln{T}}{\sqrt{B}k_{1}}+U\ln{K}\right).$
For any competitor $f$ satisfying $\|f\|_{\mathcal{H}_{i}}<1$, it must be $L_{T}(f)=O(T)$. The upper bound is trivial. To avoid this issue, we must set $U>1$ and exclude such competitors. Theorem 1 reveals how the regret bound increases with $K$, $\mathcal{R}$ and $\mathcal{A}_{T,\kappa_{i}}$. The larger the memory budget is, the smaller the regret bound will be. Theorem 1 gives an answer to Q1. If the optimal kernel $\kappa^{\ast}\in\mathcal{K}$ matches well with the examples, i.e., $\mathcal{A}_{T,\kappa^{\ast}}=o(T)$, then it is possible to achieve a $o(T)$ regret bound in the case of $\mathcal{R}=\Theta(\ln{T})$. Such a result gives an answer to Q2. According to Definition 2, learning is possible in $\mathbb{H}_{\kappa^{\ast}}$. The information-theoretical cost that the optimal kernel is unknown is $O(\sqrt{UL_{T}(f)\ln{K}})$, which is a lower order term.
Our regret bound can also derive the state-of-the-art results. Let $B=\Theta(\mathcal{A}_{T,\kappa_{i}})$. Then we obtain a $O(U\sqrt{K\mathcal{A}_{T,\kappa_{i}}})$ expected bound. Previous work proved a $O(U\sqrt{K}T^{\frac{1}{4}}\mathcal{A}^{\frac{1}{4}}_{T,\kappa_{i}})$ high-probability bound [3]. If $\mathcal{A}_{T,\kappa_{i}}=O(T)$, then we obtain a $O\left(\frac{U\sqrt{K}T}{\sqrt{B}}\right)$ expected bound which matches the upper bound in [4]. If $B=\Theta(T)$, then our regret bound matches the upper bounds in [1, 2, 10].
Next we show an algorithm-dependent bound which proves that removing a half of examples can be better than the restart technique. Let $\mathcal{J}_{i}=\{t\in[T]:|S_{i}|=\alpha\mathcal{R},b_{t,i}=1\}$ and
$\Lambda_{i}=\sum_{t\in\mathcal{J}_{i}}\left[\left\|\bar{f}^{\prime}_{t-1,i}(1)-f\right\|^{2}_{\mathcal{H}_{i}}-\left\|f^{\prime}_{t-1,i}-f\right\|^{2}_{\mathcal{H}_{i}}\right].$
There must be a constant $\xi_{i}\in(0,4]$ such that $\Lambda_{i}=\xi_{i}U^{2}|\mathcal{J}_{i}|$. Recalling that $\bar{f}^{\prime}_{t-1,i}(1)$ is the initial hypothesis after a removing operation. It is natural that if $(\bar{f}^{\prime}_{t-1,i}(1)-f)$ is close to $(f^{\prime}_{t-1,i}-f)$, then $\xi_{i}\ll 4$. The restart technique sets $\bar{f}^{\prime}_{t-1,i}(1)=0$ which implies $\Lambda_{i}\leq U^{2}|\mathcal{J}_{i}|$. In the worst case, our approach is slightly worse than restart. If $\xi_{i}$ is small enough, then our approach is much better than restart.
Theorem 2 (Algorithm-dependent Bound).
Let the conditions in Theorem 1 be satisfied. For each $i\in[K]$, let $\lambda_{i}=\frac{\sqrt{2}U}{\sqrt{5\tilde{\mathcal{A}}_{T,\kappa_{i}}}}$. If $\xi_{i}\cdot|\mathcal{J}_{i}|\leq 1$, then the expected regret of M-OMD-H satisfies,
$\forall\kappa_{i}\in\mathcal{K},f\in\mathbb{H}_{i},\quad\mathbb{E}\left[\mathrm{Reg}(f)\right]=O\left(\sqrt{L_{T}(f)\ln{K}}+\frac{UK}{B}\mathcal{A}_{T,\kappa_{i}}\ln{T}+U\sqrt{\mathcal{A}_{T,\kappa_{i}}\ln{T}}+U\ln{K}\right).$
The dominate term in Theorem 2 is $\tilde{O}\left(\frac{UK\mathcal{A}_{T,\kappa_{i}}}{B}\right)$, while the dominate term in Theorem 1 is $\tilde{O}\left(\frac{UK\mathcal{A}_{T,\kappa_{i}}}{\sqrt{B}}\right)$. The restart technique can only enjoys a regret bound given in Theorem 1. Thus removing a half of examples can be better than restart.
4.2 Smooth Loss Functions
We first define the smooth loss functions.
Assumption 3.
[19] Let $G_{1}$ and $G_{2}$ be positive constants. For any $u$ and $y$, $\ell(u,y)$ satisfies (i) $|\ell^{\prime}(u,y)|\leq G_{1}$, (ii) $|\ell^{\prime}(u,y)|\leq G_{2}\ell(u,y)$, where $\ell^{\prime}(u,y)=\frac{\mathrm{d}\,\ell(u,y)}{\mathrm{d}\,u}$.
The logistic loss function satisfies Assumption 3 with $G_{2}=1$. For the hinge loss function, M-OMD-H uniformly allocates the memory budget and induces a $O(\sqrt{K})$ factor in the regret bound. For smooth loss functions, we will propose a memory sharing scheme that can avoid the $O(\sqrt{K})$ factor.
Recalling that the memory is used to store the examples and coefficients. The main challenge is how to share the examples. To this end, we only keep a single buffer for $\{f_{t,i}\}^{K}_{i=1}$. We first rewrite $f_{t}({\bm{x}}_{t})$ as follows,
$f_{t}({\bm{x}}_{t})=\sum^{K}_{i=1}p_{t,i}f_{t,i}({\bm{x}}_{t})=\sum^{K}_{i=1}p_{t,i}\sum_{{\bm{x}}_{\tau}\in S_{i}}a_{\tau,i}\kappa_{i}({\bm{x}}_{\tau},{\bm{x}}_{t}).$
Thus we just need to ensure $S_{i}=S$ for all $i\in[K]$. For each $f_{t,i}$, we define a surrogate gradient $\nabla_{t,i}=\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\kappa_{i}({\bm{x}}_{t},\cdot)$. Note that we use the first derivative $\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})$, but not $\ell^{\prime}(f_{t,i}({\bm{x}}_{t}),y_{t})$. Let ^$\hat{\nabla}_{t,i}=0$. We can merge (4) and (5) and give a single update rule
$f_{t+1,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\langle f,\nabla_{t,i}\rangle+\mathcal{B}_{\psi_{i}}(f,f_{t,i})\right\}.$
To ensure $S_{i}=S$, we must change the sampling scheme (6). Let $b_{t,i}=b_{t}$ and $\mathrm{con}(a(i))=\mathrm{con}(a)$ for all $i\in[K]$. We define $\nu=1$ and
$Z_{t}=$ $\|\nabla_{t,i}\|_{\mathcal{H}_{i}}+G_{1}\sqrt{\kappa_{i}({\bm{x}}_{t},{\bm{x}}_{t})},$
$\mathrm{con}(a):=$ $\max_{i\in[K]}\|\kappa_{i}({\bm{x}}_{i(s_{t})},\cdot)-\kappa_{i}({\bm{x}}_{t},\cdot)\|_{\mathcal{H}_{i}}\leq\gamma_{t}.$
If $\mathrm{con}(a)$, then we replace (7) with (13).
$f_{t+1,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\langle f,\nabla_{i(s_{t}),i}\rangle+\mathcal{B}_{\psi_{i}}(f,f_{t,i})\right\},$ (13)
where $\nabla_{i(s_{t}),i}=\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\kappa_{i}({\bm{x}}_{i(s_{t})},\cdot)$. Otherwise, we replace (8) with (14).
$f_{t+1,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\left\langle f,\tilde{\nabla}_{t,i}\right\rangle+\mathcal{B}_{\psi_{i}}(f,f_{t,i})\right\}.$ (14)
If $|S|=\alpha\mathcal{R}$ and $b_{t}=1$, then let $o_{t}=1$. We remove $f_{t,i}(1)$ from $f_{t,i}$ and redefine (9) as follows
$f_{t+1,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\left\langle f,\tilde{\nabla}_{t,i}\right\rangle+\mathcal{B}_{\psi_{i}}\left(f,\bar{f}_{t,i}(2)\right)\right\}.$ (15)
Let ${\bm{p}}_{1}$ be the uniform distribution and ${\bm{p}}_{t+1}$ follow (10), in which $c_{t,i}$ follows (16).
Definition 3.
At each round $t$, $\forall\kappa_{i}\in\mathcal{K}$, let
$c_{t,i}=\left\{\begin{array}[]{ll}\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\cdot\left(f_{t,i}({\bm{x}}_{t})-\min_{j\in[K]}f_{t,j}({\bm{x}}_{t})\right)&~{}\mathrm{if}~{}\ell^{\prime}(f_{t})>0,\\ \ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\cdot\left(f_{t,i}({\bm{x}}_{t})-\max_{j\in[K]}f_{t,j}({\bm{x}}_{t})\right)&~{}\mathrm{otherwise}.\end{array}\right.$ (16)
We name this algorithm M-OMD-S (Memory bounded OMD for Smooth loss function) and present the pseudo-code in Algorithm 2.
Algorithm 2 M-OMD-S
0:  $\{\lambda_{i}\}^{K}_{i=1}$, $\{\eta_{t}\}^{T}_{t=1}$, $\mathcal{R}$, $U$.
0:  $S=\emptyset$, $f_{1,i}=0,i\in[K]$.
1:  for $t=1,2,\ldots,T$ do
2:     Receive ${\bm{x}}_{t}$
3:     for $i=1,\ldots,K$ do
4:        Find ${\bm{x}}_{i(s_{t})}$
5:        Compute $f_{t,i}({\bm{x}}_{t})$
6:     end for
7:     Output $f_{t}({\bm{x}}_{t})$ or $\hat{y}_{t}=\mathrm{sign}(f_{t}({\bm{x}}_{t}))$ and receive $y_{t}$
8:     if $\mathrm{con}(a)$ then
9:        for $i=1,\ldots,K$ do
10:           Compute $f_{t+1,i}$ following (13)
11:        end for
12:     else
13:        Sample $b_{t}\sim\mathrm{Ber}(\mathbb{P}[b_{t}=1],1)$
14:        if $b_{t}=1$ and $|S|\leq\alpha\mathcal{R}-1$ then
15:           $\forall i\in[K]$, update $f_{t+1,i}$ following (14)
16:        end if
17:        if $b_{t}=1$ $\mathrm{and}$ $|S|=\alpha\mathcal{R}$ then
18:           Remove the first $\frac{1}{2}\alpha\mathcal{R}$ support vectors from $S$
19:           $\forall i\in[K]$, update $f_{t+1,i}$ following (15)
20:           Update $S=S\cup\{({\bm{x}}_{t},y_{t})\}$
21:        end if
22:     end if
23:     Compute ${\bm{p}}_{t+1}$ by (10)
24:  end for
Lemma 2.
Let $\delta\in(0,1)$, $\alpha\mathcal{R}:=B\geq 21\ln\frac{1}{\delta}$ and $\ell$ satisfy Assumption 3. For any $\mathcal{I}_{T}$, with probability at least $1-\frac{T}{B}\Theta(\lceil\ln{T}\rceil)\delta$, the times that M-OMD-S executes removing operation on $S$ are $\left\lceil\frac{4G_{2}\hat{L}_{1:T}}{(B-\frac{4}{3}\ln\frac{1}{\delta})G_{1}}\right\rceil$ at most, in which $\hat{L}_{1:T}=\sum^{T}_{t=1}\ell(f_{t}({\bm{x}}_{t}),y_{t})$.
It is important that the times of removing operation depend on the cumulative losses of the algorithm, not the time horizon $T$.
Theorem 3.
Let $K\leq d$, $\eta_{t}$ follow (10),
$\gamma_{t}=$ $\frac{\sqrt{2\ln{K}}}{\sqrt{1+\sum^{t}_{\tau=1}|\ell^{\prime}(f_{\tau}({\bm{x}}_{\tau}),y_{\tau})|}},$ (17)
$U\leq$ $\frac{1}{8G_{2}}\sqrt{B-\frac{4}{3}\ln\frac{1}{\delta}},$
and $\lambda_{i}=\frac{2U}{G_{1}\sqrt{B}}$. Let $\mathcal{R}$ and $\ell$ satisfy Lemma 2. With probability at least $1-\frac{3T}{\alpha\mathcal{R}}\Theta(\lceil\ln{T}\rceil)\delta$, the regret of M-OMD-S satisfies
$\forall\kappa_{i}\in\mathcal{K},~{}\forall f$ $\in\mathbb{H}_{i},\quad\mathrm{Reg}(f)\leq$
$\frac{24UG_{2}L_{T}(f)}{\sqrt{B-\frac{4}{3}\ln\frac{1}{\delta}}}+UG_{1}\sqrt{B}+\frac{100U^{2}G_{2}G_{1}\ln\frac{1}{\delta}}{(1-\gamma)^{2}}+10U(1-\gamma)^{-\frac{3}{2}}\sqrt{G_{2}G_{1}\ln\frac{1}{\delta}}\sqrt{L_{T}(f)+\frac{UG_{1}\sqrt{B}}{4}},$
where $\gamma={6UG_{2}}/{\sqrt{B-\frac{4}{3}\ln\frac{1}{\delta}}}$.
In the case of $B\ll T$, the dominant term in the upper bound is $O\left(\frac{L_{T}(f)}{\sqrt{B}}\right)$, Previous work proved a $\Omega(T/B)$ lower bound [19]. Thus our result proves a new trade-off between memory constraint and regret bound which gives an answer to Q1. If $\min_{f\in\mathbb{H}_{\kappa^{\ast}}}L_{T}(f)=o(T)$, then a $\Theta(\ln{T})$ memory budget is enough to achieve a sub-linear regret which answers Q2. According to Definition 2, we also conclude that $\mathbb{H}_{\kappa^{\ast}}$ is online learnable. The penalty term for not knowing the optimal kernel is $O(U\sqrt{L_{T}(f)\ln{K}})$ which only depends on $O(\sqrt{\ln{K}})$. In this case, online kernel selection is not much harder than online kernel learning.
Theorem 3 can also recover some state-of-the-art results. Let $B=\Theta(T)$. We obtain a $O(U\sqrt{T\ln{K}})$ regret bound which is same with the regret bound in [1, 2, 10]. If $K=1$ and $B=\Theta(L_{T}(f))$, then we obtain a $O(U\sqrt{L_{T}(f)})$ regret bound matching the regret bound in [19]. Next we prove that M-OMD-S is optimal.
Theorem 4 (Lower Bound).
Let $\ell(u,y)=\ln(1+\exp(-yu))$ which satisfies Assumption 3 and $\kappa({\bm{x}},{\bm{v}})=\langle{\bm{x}},{\bm{v}}\rangle^{p}$ where $p>0$ is a constant. Let $U<\frac{\sqrt{3B}}{5}$ and =$\mathbb{H}=\{f\in\mathcal{H}_{\kappa}:\|f\|_{\mathcal{H}}\leq U\}$. There exists a $\mathcal{I}_{T}$ such that the regret of any online kernel algorithms storing $B$ examples satisfies
$\exists f\in\mathbb{H},\quad\mathbb{E}\left[\mathrm{Reg}(f)\right]=\Omega\left(U\cdot\frac{L_{T}(f)}{\sqrt{B}}\right).$
The expectation is w.r.t. the randomness of algorithm and environment.
Our lower bound improves the previous $\Omega(T/B)$ [19]. Next we show an algorithm-dependent bound. Let
$\mathcal{J}=$$\left\{t\in[T]:|S|=\alpha\mathcal{R},b_{t}=1\right\},$
$\Lambda_{i}=$ $\sum_{t\in\mathcal{J}}\left[\left\|f_{t,i}(2)-f\right\|^{2}_{\mathcal{H}_{i}}-\left\|f_{t,i}-f\right\|^{2}_{\mathcal{H}_{i}}\right].$
There must be a constant $\xi_{i}\in(0,4]$ such that $\Lambda_{i}=\xi_{i}U^{2}|\mathcal{J}|$. It is natural that if $(f_{t,i}(2)-f)$ is close to $(f_{t,i}-f)$, then $\xi_{i}\ll 4$. The restart technique sets $f_{t,i}(2)=0$ which implies $\Lambda_{i}\leq U^{2}|\mathcal{J}|$.
Theorem 5 (Algorithm-dependent Bound).
Let the conditions in Theorem 3 be satisfied and $U\leq\frac{B-\frac{4}{3}\ln\frac{1}{\delta}}{32G_{2}}$. For a fixed $i\in[K]$, if $\xi_{i}<\frac{1}{|\mathcal{J}|}$, then with probability at least $1-\frac{3T}{\alpha\mathcal{R}}\Theta(\lceil\ln{T}\rceil)\delta$, the regret of M-OMD-S satisfies,
$\forall f\in\mathbb{H}_{i},\mathrm{Reg}(f)\leq$ $\frac{32UG_{2}L_{T}(f)}{B-\frac{4}{3}\ln\frac{1}{\delta}}+\frac{144U^{2}G_{2}G_{1}\ln\frac{1}{\delta}}{(1-\gamma)^{2}}+12U(1-\gamma)^{-\frac{3}{2}}\sqrt{G_{2}G_{1}L_{T}(f)\ln\frac{1}{\delta}},$
where $\gamma=\frac{16UG_{2}}{B-\frac{4}{3}\ln\frac{1}{\delta}}$.
The dominate term in Theorem 5 is $O(\frac{UG_{2}L_{T}(f)}{B})$, while the dominate term in Theorem 3 is $O(\frac{UG_{2}L_{T}(f)}{\sqrt{B}})$. The restart technique can only enjoys a regret bound given in Theorem 3. Thus removing a half of examples can be better than restart.
5 Experiments
In this section, we will verify the following two goals,
  1. G1
    Learning is possible in the optimal hypothesis space within a small memory budget.
    We will verify the data-dependent complexities are smaller than the worst-case complexity, i.e., $\min_{\kappa_{i}}\mathcal{A}_{T,\kappa_{i}}\ll T$ and $\min_{f\in\mathbb{H}_{i},i\in[K]}L_{T}(f)\ll T$.
  2. G2
    Our algorithms perform better than baseline algorothms within a same memory budget.
5.1 Experimental Setting
We implement M-OMD-S with the logistic loss function. Thus both M-OMD-H and M-OMD-S are suitable for online binary classification tasks. We adopt the Gaussian kernel $\kappa({\bm{x}},{\bm{v}})=\exp(-\frac{\|{\bm{x}}-{\bm{v}}\|^{2}_{2}}{2\sigma^{2}})$, and choose 8 classification datasets from UCI machine learning repository  [^2] http://archive.ics.uci.edu/ml/datasets.php. The information of the datasets is shown in Table 1. All algorithms are implemented in R on a Windows machine with 2.8 GHz Core(TM) i7-1165G7 CPU. We execute each experiment 10 times with random permutation of all datasets.
Table 1: Datasets Used in Experiments
Datasets #sample # Feature Datasets #sample # Feature Datasets #sample # Feature
mushrooms 8,124 112 phishing 11,055 68 magic04 19,020 10
a9a 48,842 123 w8a 49,749 300 SUSY 50,000 18
ijcnn1 141,691 22 cod-rna 271,617 8 -
We compare with two baseline algorithms.
  • Raker: online multi-kernel learning via random features [10].
  • LKMBooks: OKS with memory constraint via random adding examples [4].
Our algorithms improve the regret bounds of Raker and LKMBooks, and we expect that our algorithms also perform better. Although there are other online kernel selection algorithms, such as OKS [26], OMKC [27], OMKR [1], we do not compare with these algorithms, since they do not limit the memory budget.
We select $K=5$ Gaussian kernel function by setting $\sigma=2^{[-2:2:6]}$. For the baseline algorithms, we tune the hyper-parameters following original papers. To be specific, for LKMBooks, there are three hyper-parameters $\nu\in(0,1)$, $\lambda$ and $\eta$. Both $\lambda$ and $\eta$ are learning rates. We reset $\lambda=10\frac{\sqrt{(1+\nu)B}}{\sqrt{1-\nu}T}$ for improving performance and tune $\nu\in(\frac{1}{2},\frac{1}{3},\frac{1}{4})$. For Raker, there is a learning rate $\eta$ and a regularization parameter $\lambda$. We tune $\eta=\frac{10^{-3:1:3}}{\sqrt{T}}$ and $\lambda\in\{0.05,0.005,0.0005\}$. For M-OMD-H and M-OMD-S, we can set an aggressive value of $U=\sqrt{B}$ following Theorem 2 and Theorem 5. We tune the learning rate $\lambda_{i}=c\frac{U}{\sqrt{B}}$ where $c\in\{2,1,0.5\}$. The value of $\eta_{t}$ follows (10). The value of $\gamma_{t}$ follows (12) and (17). For M-OMD-H, we set $M=10$.
Let $\mathrm{MR}=\frac{1}{T}\sum^{T}_{t=1}\mathbb{I}_{\hat{y}_{t}\neq y_{t}}$ be the mistake rate. Table 2 and Table 3 summary the experimental results including the average MR (AMR) of 10 random running, memory budget ($B$ or $D$), $\mathcal{A}_{T,\kappa^{\ast}}=\min_{\kappa_{i}}\mathcal{A}_{T,\kappa_{i}}$ and $L^{\ast}=\min_{f\in\mathbb{H}_{i},i\in[k]}L_{T}(f)$. $D$ is the number of random features adopted by Raker and plays a similar role with $B$. For completeness, we also report the average running time which is not a critical measurement.
5.2 Experimental Results
We separately use the hinge loss function and the logistic loss function.
5.2.1 Hinge Loss Function
Table 2 shows the results. Note that M-OMD-H keeps $5$ buffers, i.e., $S_{i},i=1,\ldots,5$, $|S_{i}|\leq B$ and LKMBooks only keeps a buffer $S$, $|S|\leq B$. Thus Raker, LKMBooks and M-OMD-H with $B=100$ enjoy a same memory budget.
We first compare the AMR of all algorithms. As a whole, M-OMD-H with $B=400$ performs best on most of datasets, while M-OMD-H with $B=100$ performs worst. The reason is that M-OMD-H uniformly allocates the memory budget over all kernels which induces a $O(\sqrt{K})$ factor on the regret bound. If we increase the memory by $K$ times, then the $O(\sqrt{K})$ factor can be removed. Thus M-OMD-H with $B=400$ performs better. The experimental results do not completely coincide with G2. It is left for future work to study whether we can avoid allocating the memory budget.
Next we analyze the value of $\mathcal{A}_{T,\kappa^{\ast}}$. Denote by $\hat{\mathcal{A}}_{T,i}:=\sum_{\tau\leq T}\left\|\nabla_{\tau,i}-\hat{\nabla}_{\tau,i}\right\|^{2}_{\mathcal{H}_{i}}\cdot\mathbb{I}_{\nabla_{\tau,i}\neq 0}$. Computing $\mathcal{A}_{T,\kappa^{\ast}}$ requires time in $O(T^{2})$. For simplicity, we use $\min_{i\in[K]}\hat{\mathcal{A}}_{T,i}$ as a proxy of $\mathcal{A}_{T,\kappa^{\ast}}$. In fact, our regret bound depends on $\hat{\mathcal{A}}_{T,i}$ which satisfies $\hat{\mathcal{A}}_{T,i}=\tilde{O}(\mathcal{A}_{T,\kappa_{i}})$. To obtain a precise estimation of $\mathcal{A}_{T,\kappa_{i}}$, we again run M-OMD-H with $M=30$ and $B=400$.
We find that $\mathcal{A}_{T,\kappa^{\ast}}<T$ on all datasets. Especially, on the mushrooms, phishing, w8a datasets, we have $\mathcal{A}_{T,\kappa^{\ast}}\ll T$. In the optimal hypothesis space, a small memory budget is enough to guarantee a small regret and thus learning is possible. The experimental results verify G1.
Table 2: Experimental Results Using the Hinge Loss Function
Algorithm mushrooms, $T=8124$ phishing, $T=11055$
AMR $B|D$ Time (s) $\mathcal{A}_{T,\kappa^{\ast}}$ AMR $B|D$ Time (s) $\mathcal{A}_{T,\kappa^{\ast}}$
Raker [10] 11.97 $\pm$ 0.93 400 1.57 - 9.13 $\pm$ 0.28 400 2.06 -
LKMBooks [4] 6.04 $\pm$ 0.62 400 1.64 - 12.81 $\pm$ 0.54 400 2.21 -
M-OMD-H 8.09 $\pm$ 2.51 100 5.62 - 19.41 $\pm$ 3.78 100 5.05 -
M-OMD-H 0.69 $\pm$ 0.07 400 10.97 220 8.87 $\pm$ 0.40 400 10.74 2849
Algorithm magic04, $T=19020$ a9a, $T=48842$
AMR $B|D$ Time (s) $\mathcal{A}_{T,\kappa^{\ast}}$ AMR $B|D$ Time (s) $\mathcal{A}_{T,\kappa^{\ast}}$
Raker [10] 32.14 $\pm$ 0.43 400 3.25 - 23.47 $\pm$ 0.23 400 9.45 -
LKMBooks [4] 23.66 $\pm$ 0.51 400 1.31 - 21.29 $\pm$ 1.36 400 10.15 -
M-OMD-H 29.08 $\pm$ 3.78 100 5.50 - 24.84 $\pm$ 1.43 100 31.39 -
M-OMD-H 22.75 $\pm$ 0.57 400 6.09 10422 19.88 $\pm$ 1.44 400 49.29 19393
Algorithm w8a, $T=49749$ SUSY, $T=50000$
AMR $B|D$ Time (s) $\mathcal{A}_{T,\kappa^{\ast}}$ AMR $B|D$ Time (s) $\mathcal{A}_{T,\kappa^{\ast}}$
Raker [10] 2.98 $\pm$ 0.00 400 11.62 - 29.62 $\pm$ 0.46 400 8.88 -
LKMBooks [4] 2.97 $\pm$ 0.00 400 23.52 - 29.37 $\pm$ 0.87 400 5.47 -
M-OMD-H 2.99 $\pm$ 0.10 100 43.20 - 36.21 $\pm$ 2.08 100 16.29 -
M-OMD-H 2.65 $\pm$ 0.12 400 84.80 4200 27.76 $\pm$ 1.04 400 29.45 28438
Algorithm ijcnn1, $T=141691$ con-rna, $T=271617$
AMR $B|D$ Time (s) $\mathcal{A}_{T,\kappa^{\ast}}$ AMR $B|D$ Time (s) $\mathcal{A}_{T,\kappa^{\ast}}$
Raker [10] 9.49 $\pm$ 0.08 400 23.54 - 11.90 $\pm$ 0.31 400 53.32 -
LKMBooks [4] 9.58 $\pm$ 0.01 400 18.85 - 12.57 $\pm$ 0.26 400 19.17 -
M-OMD-H 9.98 $\pm$ 0.29 100 33.15 - 14.59 $\pm$ 1.47 100 70.33 -
M-OMD-H 9.57 $\pm$ 0.43 400 54.81 33626 12.46 $\pm$ 0.54 400 130.09 62259
5.2.2 Logistic Loss Function
Table 3 summaries the results. As a whole, M-OMD-S performs best on most of datasets. On the phishing, SUSY and cod-rna datasets, Raker performs better than M-OMD-S. We find that the current regret analysis of Raker is not tight, and can be improved using Assumption 3. Thus it is reasonable that Raker performs similar with M-OMD-S. Besides, M-OMD-S performs better than LKMBooks on all datasets. The experimental results coincide with the theoretical analysis that is, M-OMD-S enjoys a better regret bound than LKMBooks. The experimental results verify G2.
Next we analyze the value of $L^{\ast}$. For simplicity, we use the cumulative losses of M-OMD-S, $\hat{L}_{1:T}$, as a proxy of $L^{\ast}$. Note that $L^{\ast}\leq\hat{L}_{1:T}$. Table 3 shows that $L^{\ast}<T$ on all datasets. Besides, on the mushrooms, phishing and w8a datasets, $L^{\ast}\ll T$. In the optimal hypothesis space, a small memory budget is enough to guarantee a small regret and thus learning is possible. The experimental results verify G1.
Table 3: Experimental Results Using the Logistic Loss Function
Algorithm mushrooms, $T=8124$ phishing, $T=11055$
AMR $B|D$ Time (s) $L^{\ast}$ AMR $B|D$ Time (s) $L^{\ast}$
Raker [10] 11.47 $\pm$ 0.75 400 1.64 - 9.07 $\pm$ 0.44 400 2.05 -
LKMBooks [4] 10.01 $\pm$ 1.44 400 2.12 - 15.86 $\pm$ 1.28 400 1.94 -
M-OMD-S 2.70 $\pm$ 0.32 400 1.74 160 10.38 $\pm$ 0.31 400 3.55 3012
Algorithm magic04, $T=19020$ a9a, $T=48842$
AMR $B|D$ Time (s) $L^{\ast}$ AMR $B|D$ Time (s) $L^{\ast}$
Raker [10] 30.74 $\pm$ 1.11 400 3.56 - 23.13 $\pm$ 0.17 400 9.90 -
LKMBooks [4] 25.72 $\pm$ 0.64 400 1.30 - 22.38 $\pm$ 1.29 400 10.98 -
M-OMD-S 23.97 $\pm$ 0.23 400 1.81 10160 19.43 $\pm$ 0.16 400 10.41 20194
Algorithm w8a, $T=49749$ SUSY, $T=50000$
AMR $B|D$ Time (s) $L^{\ast}$ AMR $B|D$ Time (s) $L^{\ast}$
Raker [10] 2.97 $\pm$ 0.00 400 12.06 - 28.62 $\pm$ 0.44 400 9.02 -
LKMBooks [4] 2.98 $\pm$ 0.01 400 6.10 - 30.72 $\pm$ 1.33 400 6.45 0
M-OMD-S 2.83 $\pm$ 0.03 400 21.44 5116 29.03 $\pm$ 0.16 400 8.24 28582
Algorithm ijcnn1, $T=141691$ con-rna, $T=271617$
AMR $B|D$ Time (s) $L^{\ast}$ AMR $B|D$ Time (s) $L^{\ast}$
Raker [10] 9.53 $\pm$ 0.03 400 24.94 - 12.27 $\pm$ 0.57 400 52.24 -
LKMBooks [4] 9.57 $\pm$ 0.00 400 16.95 - 13.91 $\pm$ 1.31 400 25.55 -
M-OMD-S 9.42 $\pm$ 0.05 400 19.33 39208 13.09 $\pm$ 0.04 400 31.20 94326
6 Conclusion
Learnability is an essential problem in online kernel selection with memory constraint. Previous work gave a negative result on learnability via worst-case regret analysis. We give positive results via data-dependent regret analysis. We characterize the regret bounds via two data-dependent complexities, i.e., kernel alignment and the cumulative losses of competitive hypothesis, and explicitly balance the data-dependent regret bounds and memory constraint. If kernel functions match well with the data, that is, the kernel alignment and the cumulative losses are sub-linear, then sub-linear regret can be achieved within a $\Theta(\ln{T})$ memory constraint and the corresponding hypothesis space is learnable.
Compared with the worst-case regret analysis which uses data size as data complexity, data-dependent regret analysis uses tighter data complexities to characterize the hardness of problems, and gives more insightful relationship among online learnability, memory constraint and data complexity. For the future work, it is interesting to study new data complexities under different loss functions.
Deploying machine learning models on computational devices with limited computational resources is necessary for practical applications. Our algorithms limit the memory budget to $O(\ln{T})$, and can naturally be deployed on computational devices with limited memory. Even if the deployed machine learning models are not kernel classifiers, our work can guide how to allocate the memory resources properly. For instance, (i) our regret bounds show the optimal values of memory budget depend on the data-dependent complexities; (ii) our second algorithm provides the idea of memory sharing.